# Multi-fidelity Bayesian Optimization: A Review

### v2

@@ 7 Further research topics
Recent advances in MFO and especially in BO have focused on solving intricate yet important design optimization problems, which stem from the nature of problems or from the inherent limitations of the generic BO. In this section, we briefly describe sophisticated BO and several MFO approaches to addressing some of these problems, including constrained optimization (Section 7.1), high-dimensional optimization (Section 7.2), optimization under uncertainty (Section 7.3), and multi-objective optimization (Section 7.4). By doing so, we expect to shed light on potential opportunities for future research in MF BO.
@@ 7.1 Constrained optimization
There exist three common BO-based approaches to solving problem (59). The first approach formulates constrained acquisition functions $\alpha_{\text{c}}({\bf x})$ that consider the influence of constraints on our selection of new design points. The second approach reformulates problem (59) as an unconstrained problem that can be addressed by the generic BO. The third approach employs TS, as described in Algorithm 2, via realizations of GP posteriors (i.e., sample paths) for the objective and constraint functions.
@@
The augmented Lagrangian method starts with initial values of $\rho_{0}$ and $\boldsymbol{\lambda}$. A sufficiently large initial value of $\rho_{0}$ is required to enforce feasibility. It then sequentially finds the minimizer of the augmented Lagrangian corresponding to the given values of $\boldsymbol{\lambda}$ and $\rho_{0}$, and updates them for the next iteration with the resulting minimizer. Algorithm 6 implements the augmented Lagrangian method (not in the BO setting), where $\eta_{1}$ and $\eta_{2}$ in 1 are two small thresholds for checking the termination conditions in 6.
In the context of BO, the idea is to construct in 3 of Algorithm 6 a GP for the augmented Lagrangian $\mathcal{L}({\bf x}|\boldsymbol{\lambda}^{k-1},\rho^{k-1}_{0})$ using observations of several design variable vectors and the corresponding values of the objective and constraint functions. Based on this GP, the next design point ${\bf x}^{k}$ in 4 is found by maximizing EI, and the new parameters $\rho_{0}$ and $\boldsymbol{\lambda}$ are updated accordingly. The advantage of this approach is that it, via BO, can increase the chance of finding a global solution as compared with the traditional augmented Lagrangian method, which is a local optimization algorithm.
@@
The third approach to solving problem (59) is via TS [262]. Let $q^{k}_{\text{H},i}({\bf x})$ and $h^{k}_{\text{H},i}({\bf x})$ denote the realizations (i.e., sample paths) of $\hat{f}_{\text{H}}({\bf x})$ and $\hat{g}_{\text{H},i}({\bf x})$ at iteration $k$ of TS, respectively. Assume that we perform 10 of Algorithm 2 in a batch setting of TS and obtain a total of $q$ new candidates, denoted by ${\bf x}^{k}_{1},\dots,{\bf x}^{k}_{q}$. Let $\mathcal{F}^{k}=\{{\bf x}^{k}_{l}|\,h^{k}_{\text{H},i}({\bf x})\leq 0,\forall l\in\{1,\dots q\},\forall i\in\{1,\dots I\}\}$ be a set of points whose realizations are feasible. If $\mathcal{F}^{k}\neq\emptyset$, we select one of its elements that minimizes $q^{k}_{\text{H},i}({\bf x})$. Otherwise, we select a candidate with a minimum value of the total violation $\sum_{i=1}^{I}\max\{0,h_{\text{H},i}({\bf x})\}$.
@@ 7.2 High-dimensional optimization
  • Exploiting low, active/effective dimensional subspace of the design variables (Section 7.2.1).
  • Exploiting the additive structures of the objective and/or constraint functions (Section 7.2.2).
  • Trust-region BO (Section 7.2.3).
@@ 7.2.2 Additive structure approach
Using the assumption in Eq. 65, we can construct component GP models for $f_{m}({\bf x}_{m})$ and maximize the acquisition functions formulated from these models to progress BO, regardless of our knowledge on the structure of $f_{m}({\bf x}_{m})$. If ${\bf x}^{k}_{m}$ is the new point from maximizing the acquisition function $m$, then the new design point is ${\bf x}^{k}=\bigcup_{m=1}^{M}{\bf x}^{k}_{m}$. Alternatively, ${\bf x}^{k}$ can be found by maximizing an acquisition function derived from an overall GP for $f_{\text{H}}({\bf x})$ whose mean (or covariance) function is assigned as the sum of mean (or covariance) functions of the component GPs.
@@
Different combinations of HF and LF data may completely change the additive form in Eq. 65 because learning possible additive structures is data-driven. Thus, examining the sensitivity of decomposition structures with respect to the data used may be the first step if one wishes to use the additive structure approach for MF BO.
@@ 7.2.3 Trust-region BO (TuRBO) approach
Assume that TuRBO maintains $m$ trust regions, i.e., $\text{TR}_{1},\dots,\text{TR}_{m}$, at its iteration $k$. The local GPs constructed from these trust regions are denoted by $\mathcal{GP}_{1}^{k},\dots,\mathcal{GP}_{m}^{k}$. In each iteration, TuRBO selects a batch of $q$ new design points, i.e., $\{{\bf x}_{1}^{k},\dots,{\bf x}_{q}^{k}\}$, which are drawn from the union of the trust regions, and then updates the local GPs associated with the trust regions from which the new design points are drawn [277]. This can be done via performing TS described in Algorithm 2.
@@
  • Use local GP-based MF surrogates for local BO runs and then adopt the no-fidelity consideration approach (Section 5.2.1) to update these surrogates after ${\bf x}_{i}^{k}$ is obtained by solving problem (66).
  • Use local GP-based MF surrogates for local BO runs and then adopt the sequential selection approach (Section 5.2.3) to select fidelity variables after ${\bf x}_{i}^{k}$ is obtained by solving problem (66). In this way, the fidelity-query function $\gamma({\bf x}_{i}^{k},t)$ described in Eq. 57 can be formulated for the trust region from which ${\bf x}_{i}^{k}$ is drawn.
@@ 7.3 Optimization under uncertainty
Although there is a rich literature on techniques for optimization in the face of uncertainty, processing this class of optimization is still a challenging task in engineering design. In this section, we briefly review recent applications of MF modeling methods and BO in two branches of optimization under uncertainty: robust optimization (RO) (Section 7.3.1) and reliability-based optimization (RBO) (Section 7.3.2). We also attempt to provide possible directions for future research on MF BO to address these problems.
Let ${\bf s}$ denote a vector of random parameters that encapsulates uncertainty in our problem. If a design variable is random, then it can be defined as a sum of a nominal variable, which is considered a design variable, and a random parameter before the optimization problem is formulated. We describe uncertainty in ${\bf s}$ using an interval $[{\bf s}_{\text{l}},{\bf s}_{\text{u}}]$ for set-based uncertainty and a PDF $p({\bf s})$ for probabilistic uncertainty. The latter is the focal point of our discussion below.
@@ 7.3.1 Robust optimization
An RO problem is often formulated using one of the following three concepts: absolute robustness, relative robustness, and less variance [281]. The absolute robustness formulates the RO problem based on the worst values of the objective and constraint functions under set-based uncertainty with a fixed set $[{\bf s}_{\text{l}},{\bf s}_{\text{u}}]$. This problem is to minimize the worst value of the objective function under the constraints on the worst values of constraint functions, which is also called the worst-case scenario or minimax approach [see e.g., 282, 283]. The relative robustness formulates the problem under set-based uncertainty to maximize the gap between the nominal value and the worst-case value of the objective function. The less variance concept formulates the problem based on probabilistic uncertainty.
@@
In the less variance concept, a design that is less sensitive to uncertainty is considered more robust. Thus, we wish to minimize the mean and variance of the objective function simultaneously, thereby formulating a bi-objective RO problem which will be discussed in Section 7.4. Nevertheless, we can reformulate this bi-objective RO problem using the following weighted-sum formulation:
@@ 7.3.2 Reliability-based optimization
Although problem (68) is very important for optimization under uncertainty, it is too difficult to solve. The difficulty arises from three aspects. First, checking the feasibility of a candidate solution is nontrivial because the computation of probabilistic constraints is generally an NP-hard problem [299]. Second, finding an exact optimal solution may be impossible because the feasible region defined by the probabilistic constraints is generally non-convex [300]. Third, estimating the mean of the objective is computationally demanding, which is discussed in Section 7.3.1.
@@
In engineering design optimization, problem (68) is often solved approximately. Classical RBO approximation techniques fall into three distinct categories: top-level, mono-level, and decoupled approaches. A benchmark study and a comprehensive review of methods in each of these approaches can be found in Aoues and Chateauneuf [303] and Valdebenito and Schuëller [304], respectively. Other approximation techniques solve problem (68) sequentially by leveraging surrogates for $\mathbb{P}_{\bf s}\left[g_{\text{H},i}({\bf x},{\bf s})\leq 0\right]$ or $g_{\text{H},i}({\bf x},{\bf s})$. Figure 14 shows different schemes of the surrogate-based RBO, which is discussed in Moustapha and Sudret [302].
@@
Although many efficient methods, inspired by Bichon et al. [306], have adopted MF BO to facilitate reliability analyses  [see e.g., 310, 311, 312, 313], the use of MF BO for addressing RBO problems is still an open issue. This may be attributed to the conflict between the accuracy and efficiency properties of a good optimization algorithm, as mentioned in Section 1. While we focus on improving the accuracy by using sophisticated MF reliability analysis methods, we unintentionally require excessive computer time or storage, thereby reducing the optimization efficiency.
The above observations present potential opportunities for MF BO to address problem (68). In addition to improving reliability analyses, considerations should include other important aspects such as improving the solution quality, evaluating the feasibility, selecting fidelities, and addressing uncertainty due to the use of LF models. We believe that the sequential selection approach (Section 5.2) would be suitable for such serious considerations because it can decouple the selections of design variables $\bf x$, random parameters $\bf s$, and fidelity variables for different purposes. After these considerations, MF BO may look on a far horizon to explore its possibility of solving RBO problems with joint probabilistic constraints, which are also very important for decision-making under uncertainty, but generally more difficult to solve than problem (68).
@@ Acknowledgments
Let ${\boldsymbol{\phi}}=[\boldsymbol{\beta}_{\text{m}}^{\intercal},\sigma^{2}_{\text{y}},\boldsymbol{\phi}_{\text{x}}^{\intercal}]^{\intercal}$ denote the vector of hyperparameters of the GP model. We can find the optimal value ${\boldsymbol{\phi}^{\star}}$ of ${\boldsymbol{\phi}}$ by maximizing the likelihood of ${\bf Y}({\bf X})$ model derived from Eq. A.4.
@@
Let ${\bf F}({\bf X})=[{\bf f}^{\intercal}({\bf x}^{1}),\dots,{\bf f}^{\intercal}({\bf x}^{N})]^{\intercal}\in\mathbb{R}^{TN}$ be a vector that concatenates the vectors ${\bf f}({\bf x}^{i})$, $i=1,\dots,N$. Let ${\bf M}({\bf X})=[{\bf m}^{\intercal}({\bf x}^{1}|\boldsymbol{\beta}_{\text{m}}),\dots,{\bf m}^{\intercal}({\bf x}^{N}|\boldsymbol{\beta}_{\text{m}})]^{\intercal}\in\mathbb{R}^{TN}$ be a vector that concatenates the mean vectors ${\bf m}({\bf x}^{i})$. According to the GP prior in Eq. B.1, the distribution of ${\bf F}({\bf X})$ reads
@@
By maximizing the likelihood of ${\bf Y}({\bf X})$ model in Eq. B.3, we obtain the optimal set of hyperparameters $\boldsymbol{\phi}=\{\boldsymbol{\phi}_{\text{x}},\boldsymbol{\Sigma}_{\text{y}},\boldsymbol{\beta}_{\text{m}}\}$. Then, the posterior predictive PDF at $\overline{\bf x}$ can be derived as
@@
As described in Section 6.2, the shapes of NACA 4-digit airfoils are parameterized by three parameters: $c_{\max}$–the height measured from the chord line to the point where the mean camber line has the largest curvature, $x_{\max}$–the position of $c_{\max}$ in horizontal coordinates, and $t_{\max}$–the maximum thickness of the airfoil, see Fig. 10(a) [248]. The airfoil sections are created from the mean camber line and a thickness distribution drawn perpendicular to the mean camber line.
@@
where $a_{0}=0.2969$, $a_{1}=0.126$, $a_{2}=0.3516$, $a_{3}=0.2843$, and $a_{4}=-0.1015$ for an open trailing edge or $-0.1035$ for a closed trailing edge.
@@
Let ${\bf x}=[x_{1},\dots,x_{11}]^{\intercal}=\left[r_{\text{LE}},x_{\text{up}},y_{\text{up}},\kappa_{\text{xxup}},x_{\text{lo}},y_{\text{lo}},\kappa_{\text{xxlo}},y_{\text{TE}},\Delta y_{\text{TE}},\alpha_{\text{TE}},\beta_{\text{TE}}\right]^{\intercal}$ represent a vector containing the geometrical features of a PARSEC airfoil, see Fig. 10(b) and Table 6. The vertical coordinates of the upper and lower surfaces of the PARSEC airfoil section are obtained from the following polynomial equations [252]:

### v3

@@ 7 Further research topics
Recent advances in MFO and especially in BO have focused on solving intricate yet important design optimization problems, which stem from the nature of problems or from the inherent limitations of the generic BO. In this section, we briefly describe sophisticated BO and several MFO approaches to addressing some of these problems, including constrained optimization (Section˜7.1), high-dimensional optimization (Section˜7.2), optimization under uncertainty (Section˜7.3), and multi-objective optimization (Section˜7.4). By doing so, we expect to shed light on potential opportunities for future research in MF BO.
@@ 7.1 Constrained optimization
There exist three common BO-based approaches to solving problem (59). The first approach formulates constrained acquisition functions $\alpha_{\text{c}}({\bf x})$ that consider the influence of constraints on our selection of new design points. The second approach reformulates problem (59) as an unconstrained problem that can be addressed by the generic BO. The third approach employs TS, as described in Algorithm˜2, via realizations of GP posteriors (i.e., sample paths) for the objective and constraint functions.
@@
The augmented Lagrangian method starts with initial values of $\rho_{0}$ and $\boldsymbol{\lambda}$. A sufficiently large initial value of $\rho_{0}$ is required to enforce feasibility. It then sequentially finds the minimizer of the augmented Lagrangian corresponding to the given values of $\boldsymbol{\lambda}$ and $\rho_{0}$, and updates them for the next iteration with the resulting minimizer. Algorithm˜6 implements the augmented Lagrangian method (not in the BO setting), where $\eta_{1}$ and $\eta_{2}$ in Line 1 are two small thresholds for checking the termination conditions in Line 6.
In the context of BO, the idea is to construct in Line 3 of Algorithm˜6 a GP for the augmented Lagrangian $\mathcal{L}({\bf x}|\boldsymbol{\lambda}^{k-1},\rho^{k-1}_{0})$ using observations of several design variable vectors and the corresponding values of the objective and constraint functions. Based on this GP, the next design point ${\bf x}^{k}$ in Line 4 is found by maximizing EI, and the new parameters $\rho_{0}$ and $\boldsymbol{\lambda}$ are updated accordingly. The advantage of this approach is that it, via BO, can increase the chance of finding a global solution as compared with the traditional augmented Lagrangian method, which is a local optimization algorithm.
@@
The third approach to solving problem (59) is via TS [262]. Let $q^{k}_{\text{H},i}({\bf x})$ and $h^{k}_{\text{H},i}({\bf x})$ denote the realizations (i.e., sample paths) of $\hat{f}_{\text{H}}({\bf x})$ and $\hat{g}_{\text{H},i}({\bf x})$ at iteration $k$ of TS, respectively. Assume that we perform Line 10 of Algorithm˜2 in a batch setting of TS and obtain a total of $q$ new candidates, denoted by ${\bf x}^{k}_{1},\dots,{\bf x}^{k}_{q}$. Let $\mathcal{F}^{k}=\{{\bf x}^{k}_{l}|\,h^{k}_{\text{H},i}({\bf x})\leq 0,\forall l\in\{1,\dots q\},\forall i\in\{1,\dots I\}\}$ be a set of points whose realizations are feasible. If $\mathcal{F}^{k}\neq\emptyset$, we select one of its elements that minimizes $q^{k}_{\text{H},i}({\bf x})$. Otherwise, we select a candidate with a minimum value of the total violation $\sum_{i=1}^{I}\max\{0,h_{\text{H},i}({\bf x})\}$.
@@ 7.2 High-dimensional optimization
  • Exploiting low, active/effective dimensional subspace of the design variables (Section˜7.2.1).
  • Exploiting the additive structures of the objective and/or constraint functions (Section˜7.2.2).
  • Trust-region BO (Section˜7.2.3).
@@ 7.2.2 Additive structure approach
Using the assumption in Eq.˜65, we can construct component GP models for $f_{m}({\bf x}_{m})$ and maximize the acquisition functions formulated from these models to progress BO, regardless of our knowledge on the structure of $f_{m}({\bf x}_{m})$. If ${\bf x}^{k}_{m}$ is the new point from maximizing the acquisition function $m$, then the new design point is ${\bf x}^{k}=\bigcup_{m=1}^{M}{\bf x}^{k}_{m}$. Alternatively, ${\bf x}^{k}$ can be found by maximizing an acquisition function derived from an overall GP for $f_{\text{H}}({\bf x})$ whose mean (or covariance) function is assigned as the sum of mean (or covariance) functions of the component GPs.
@@
Different combinations of HF and LF data may completely change the additive form in Eq.˜65 because learning possible additive structures is data-driven. Thus, examining the sensitivity of decomposition structures with respect to the data used may be the first step if one wishes to use the additive structure approach for MF BO.
@@ 7.2.3 Trust-region BO (TuRBO) approach
Assume that TuRBO maintains $m$ trust regions, i.e., $\text{TR}_{1},\dots,\text{TR}_{m}$, at its iteration $k$. The local GPs constructed from these trust regions are denoted by $\mathcal{GP}_{1}^{k},\dots,\mathcal{GP}_{m}^{k}$. In each iteration, TuRBO selects a batch of $q$ new design points, i.e., $\{{\bf x}_{1}^{k},\dots,{\bf x}_{q}^{k}\}$, which are drawn from the union of the trust regions, and then updates the local GPs associated with the trust regions from which the new design points are drawn [277]. This can be done via performing TS described in Algorithm˜2.
@@
  • Use local GP-based MF surrogates for local BO runs and then adopt the no-fidelity consideration approach (Section˜5.2.1) to update these surrogates after ${\bf x}_{i}^{k}$ is obtained by solving problem (66).
  • Use local GP-based MF surrogates for local BO runs and then adopt the sequential selection approach (Section˜5.2.3) to select fidelity variables after ${\bf x}_{i}^{k}$ is obtained by solving problem (66). In this way, the fidelity-query function $\gamma({\bf x}_{i}^{k},t)$ described in Eq.˜57 can be formulated for the trust region from which ${\bf x}_{i}^{k}$ is drawn.
@@ 7.3 Optimization under uncertainty
Although there is a rich literature on techniques for optimization in the face of uncertainty, processing this class of optimization is still a challenging task in engineering design. In this section, we briefly review recent applications of MF modeling methods and BO in two branches of optimization under uncertainty: robust optimization (RO) (Section˜7.3.1) and reliability-based optimization (RBO) (Section˜7.3.2). We also attempt to provide possible directions for future research on MF BO to address these problems.
Let $\bf s$} denote a vector of random parameters that encapsulates uncertainty in our problem. If a design variable is random, then it can be defined as a sum of a nominal variable, which is considered a design variable, and a random parameter before the optimization problem is formulated. We describe uncertainty in $\bf s$} using an interval ${\bf s}_{\text{l}},{\bf s}_{\text{u}}]$ for set-based uncertainty and a PDF $p({\bf s})$ for probabilistic uncertainty. The latter is the focal point of our discussion below.
@@ 7.3.1 Robust optimization
An RO problem is often formulated using one of the following three concepts: absolute robustness, relative robustness, and less variance [281]. The absolute robustness formulates the RO problem based on the worst values of the objective and constraint functions under set-based uncertainty with a fixed set ${\bf s}_{\text{l}},{\bf s}_{\text{u}}]$. This problem is to minimize the worst value of the objective function under the constraints on the worst values of constraint functions, which is also called the worst-case scenario or minimax approach [see e.g., 282, 283]. The relative robustness formulates the problem under set-based uncertainty to maximize the gap between the nominal value and the worst-case value of the objective function. The less variance concept formulates the problem based on probabilistic uncertainty.
@@
In the less variance concept, a design that is less sensitive to uncertainty is considered more robust. Thus, we wish to minimize the mean and variance of the objective function simultaneously, thereby formulating a bi-objective RO problem which will be discussed in Section˜7.4. Nevertheless, we can reformulate this bi-objective RO problem using the following weighted-sum formulation:
@@ 7.3.2 Reliability-based optimization
Although problem (68) is very important for optimization under uncertainty, it is too difficult to solve. The difficulty arises from three aspects. First, checking the feasibility of a candidate solution is nontrivial because the computation of probabilistic constraints is generally an NP-hard problem [299]. Second, finding an exact optimal solution may be impossible because the feasible region defined by the probabilistic constraints is generally non-convex [300]. Third, estimating the mean of the objective is computationally demanding, which is discussed in Section˜7.3.1.
@@
In engineering design optimization, problem (68) is often solved approximately. Classical RBO approximation techniques fall into three distinct categories: top-level, mono-level, and decoupled approaches. A benchmark study and a comprehensive review of methods in each of these approaches can be found in Aoues and Chateauneuf [303] and Valdebenito and Schuëller [304], respectively. Other approximation techniques solve problem (68) sequentially by leveraging surrogates for $\mathbb{P}_{\bf s}\left[g_{\text{H},i}({\bf x},{\bf s})\leq 0\right]$ or $g_{\text{H},i}({\bf x},{\bf s})$. Figure˜14 shows different schemes of the surrogate-based RBO, which is discussed in Moustapha and Sudret [302].
@@
Although many efficient methods, inspired by Bichon et al. [306], have adopted MF BO to facilitate reliability analyses  [see e.g., 310, 311, 312, 313], the use of MF BO for addressing RBO problems is still an open issue. This may be attributed to the conflict between the accuracy and efficiency properties of a good optimization algorithm, as mentioned in Section˜1. While we focus on improving the accuracy by using sophisticated MF reliability analysis methods, we unintentionally require excessive computer time or storage, thereby reducing the optimization efficiency.
The above observations present potential opportunities for MF BO to address problem (68). In addition to improving reliability analyses, considerations should include other important aspects such as improving the solution quality, evaluating the feasibility, selecting fidelities, and addressing uncertainty due to the use of LF models. We believe that the sequential selection approach (Section˜5.2) would be suitable for such serious considerations because it can decouple the selections of design variables $\bf x$, random parameters $\bf s$, and fidelity variables for different purposes. After these considerations, MF BO may look on a far horizon to explore its possibility of solving RBO problems with joint probabilistic constraints, which are also very important for decision-making under uncertainty, but generally more difficult to solve than problem (68).
@@ Acknowledgments
Let ${\boldsymbol{\phi}}=[\boldsymbol{\beta}_{\text{m}}^{\intercal},\sigma^{2}_{\text{y}},\boldsymbol{\phi}_{\text{x}}^{\intercal}]^{\intercal}$ denote the vector of hyperparameters of the GP model. We can find the optimal value $\boldsymbol{\phi}^{\star}$} of $\boldsymbol{\phi}$} by maximizing the likelihood of ${\bf Y}({\bf X})$ model derived from Eq.˜A.4.
@@
Let ${\bf F}({\bf X})=[{\bf f}^{\intercal}({\bf x}^{1}),\dots,{\bf f}^{\intercal}({\bf x}^{N})]^{\intercal}\in\mathbb{R}^{TN}$ be a vector that concatenates the vectors ${\bf f}({\bf x}^{i})$, $i=1,\dots,N$. Let ${\bf M}({\bf X})=[{\bf m}^{\intercal}({\bf x}^{1}|\boldsymbol{\beta}_{\text{m}}),\dots,{\bf m}^{\intercal}({\bf x}^{N}|\boldsymbol{\beta}_{\text{m}})]^{\intercal}\in\mathbb{R}^{TN}$ be a vector that concatenates the mean vectors ${\bf m}({\bf x}^{i})$. According to the GP prior in Eq.˜B.1, the distribution of ${\bf F}({\bf X})$ reads
@@
By maximizing the likelihood of ${\bf Y}({\bf X})$ model in Eq.˜B.3, we obtain the optimal set of hyperparameters $\boldsymbol{\phi}=\{\boldsymbol{\phi}_{\text{x}},\boldsymbol{\Sigma}_{\text{y}},\boldsymbol{\beta}_{\text{m}}\}$. Then, the posterior predictive PDF at $\overline{\bf x}$ can be derived as
@@
As described in Section˜6.2, the shapes of NACA 4-digit airfoils are parameterized by three parameters: $c_{\max}$–the height measured from the chord line to the point where the mean camber line has the largest curvature, $x_{\max}$–the position of $c_{\max}$ in horizontal coordinates, and $t_{\max}$–the maximum thickness of the airfoil, see Fig.˜10(a) [248]. The airfoil sections are created from the mean camber line and a thickness distribution drawn perpendicular to the mean camber line.
@@
where $a_{0}=0.2969$, $a_{1}=0.126$, $a_{2}=0.3516$, $a_{3}=0.2843$, and $a_{4}=-0.1015$ for an open trailing edge or 0.1035-0.1035 for a closed trailing edge.
@@
Let ${\bf x}=[x_{1},\dots,x_{11}]^{\intercal}=\left[r_{\text{LE}},x_{\text{up}},y_{\text{up}},\kappa_{\text{xxup}},x_{\text{lo}},y_{\text{lo}},\kappa_{\text{xxlo}},y_{\text{TE}},\Delta y_{\text{TE}},\alpha_{\text{TE}},\beta_{\text{TE}}\right]^{\intercal}$ represent a vector containing the geometrical features of a PARSEC airfoil, see Fig.˜10(b) and Table˜6. The vertical coordinates of the upper and lower surfaces of the PARSEC airfoil section are obtained from the following polynomial equations [252]: